# When Bias Pretends to Be Truth: How Spurious Correlations Undermine Hallucination Detection in LLMs

### v1

@@ 1 Introduction
Figure 1: Illustration of hallucination generation and detection failure due to spurious correlations; high-confidence hallucinations arise from biases in training data, bypassing detection and mitigation strategies.
@@ 2.2 Spurious Correlation
Spurious correlations, also called “shortcuts,” are non-causal statistical dependencies in training data and significantly contribute to hallucinations in language models. Recent studies demonstrate that such correlations amplify erroneous outputs, often with high confidence. Multimodal research, for instance, shows object hallucinations being exacerbated by misleading co-occurrences in datasets (hosseini2025seeing; hu2025causal). Similarly, purely textual models suffer from biases like attestation and frequency biases, resulting in incorrect entailments or factual assertions derived from superficial patterns (mckenna2023sources). Conceptual-level spurious correlations are widespread in both fine-tuning and in-context learning settings, posing substantial mitigation challenges across modeling paradigms (zhou2023explore; yuan2024llms). Methods like high-similarity pruning and causal interventions have been proposed to address knowledge-shortcut hallucinations (wang2025kshseek; licausally), yet their efficacy is limited to particular contexts, and their performance under strong correlations remains unclear.
@@
3 Empirical Evaluation of Hallucination Detection under Spurious Correlation
@@ 3.1 Experimental Setting
Figure 2: AUROC of different hallucination detection methods versus $\rho$. Left: Experimental results of pretrained models. Right: Experimental results of models continue pretrained from SmolLM2-1.7B. All the detection methods’ AUROC drops when $\rho$ increases, showing that spurious correlation hinders hallucination detection methods. We selected the linear probing layer that performed best on the training set.
@@
Following (AllenZhu-icml2024-tutorial; allen2024physics33), we generate profiles for 20,000 individuals, each containing six attributes: date of birth, birth city, university, major, employer, and employer city. To construct both pretraining and supervised fine-tuning datasets, we first design a diverse set of text templates for describing profiles and then embed each individual’s information into natural texts based on these templates (see Table 2 for examples). We uniformly divide the profiles into three subsets—pretraining, instruction fine-tuning, and testing—to ensure balanced representation and prevent overlap. The pretraining set includes the first 10,000 individuals, each represented by 50 diverse text templates; the fine-tuning set uses the first 5,000 of these individuals, generating 30 question–answer pairs per individual. The remaining individuals are reserved exclusively for testing and hallucination detection evaluation. A detailed description of dataset construction is provided in Appendix E.1.
Model Configuration and Training Procedure
We conduct experiments using GPT2-like models (modded_nanogpt_2024) of various sizes, detailed in Table 3. The training procedures are described in Appendix E.
@@
To systematically investigate spurious correlations, we adopt a controlled methodology: each individual’s full name is composed of a first name, middle name, and surname, each randomly selected from distinct sets without repetition. We then associate surnames with specific attributes using a probabilistic mapping that simulates realistic patterns (e.g., surnames ending in kov are often linked to Russian birthplaces). To control correlation strength, we introduce a coefficient $\rho$ (ranging from 0 to 1), representing the probability that a surname directly determines its associated attribute. Thus, with probability $\rho$, the attribute matches the surname-based mapping; otherwise, it is uniformly sampled from all possible values, independent of the surname. This approach enables precise manipulation of correlation strength to rigorously evaluate existing hallucination detection methods.
Figure 3: Performance of fine-tuned models with different sizes under varying correlation coefficients. (Left) The test accuracy for factual recall questions regarding known individuals. (Right) The refusal rate when queried about unknown individuals. Different colors correspond to different model sizes.
@@ 3.2 Results
As shown in Figure 2, although some methods perform well when $\rho=0$, their performance degrades sharply as $\rho$ increases (e.g., $\rho=0.9$), with most methods failing to maintain reasonable precision. The hallucination detection methods include perplexity, logit entropy, window entropy, attention score, and linear probing. For details of hallucination detection methods, see Table LABEL:tab:hallucination_methods.
@@
To verify whether the previously identified spurious-correlation-driven failure persists under a knowledge injection setting, we extend our investigation to real LLMs fine-tuned on synthetic datasets. Using SmolLM2-1.7B (allal2025smollm2) as the base model, we conduct continual pre-training and supervised instruction fine-tuning on an intermediate checkpoint under a data injection scenario. As shown in Figure 2, when $\rho$ is high, all evaluated methods still exhibit low precision, confirming that the same failure mode remains even at the 1.7B scale.
@@
We examine how spurious correlations impact the effectiveness of refusal fine-tuning, a method that encourages models to explicitly indicate uncertainty with phrases such as I don’t know when encountering uncertain or out-of-distribution inputs. Following  zhang2024r; cheng2024can, we add refusal examples during instruction fine-tuning. They match the structure of the fine-tuning questions but mention unseen individuals, with the target response consistently set to “I don’t know” (see Table 2).
@@
4 Validation on Real World LLM Models
@@ 4.1 Experimental Setups
We conduct experiments using both open-source models and API-based commercial models to validate our claims on spurious correlations and hallucination detection in real-world LLM settings. Specifically, we select the following models for evaluation: GPT-5 (openai2025gpt5), DeepSeek V3 (liu2024deepseek), GPT-OSS-20B (agarwal2025gpt), and Qwen-30B-A3B-Instruct (qwen2025qwen25technicalreport).
@@
For each question–answer pair $(x,y)$, we obtain a consensus model answer $f^{*}(x)$ by running the model $f$ ten times and taking a majority vote over the outputs $f_{1}(x),f_{2}(x),\dots,f_{10}(x)$. We then extract entities from both the question and the consensus answer using an entity extractor $e(\cdot)$ (by prompting LLM) and compute their co-occurrence strength using the Jaccard similarity (jaccard1908nouvelles):
@@
Intuitively, a higher Jaccard similarity indicates stronger associative priors between the entities in questions and model-generated answers, effectively serving as a proxy for larger $\rho$. We compute these Jaccard similarity scores for all samples and group them into five buckets based on their values, from $T_{1}$ (highest similarity) to $T_{5}$ (lowest). This bucketing allows us to analyze model behavior under different levels of spurious correlation.
@@ 4.2 Results
Figure 5: Hallucination detection performance versus entity co-occurrence. Left: GPT-OSS-20B. Right: Qwen-30B-A3B-Instruct. AUROC scores decrease consistently as Jaccard overlap increases, across all evaluated detection methods including perplexity, window entropy, logit entropy, attention-score heuristics, and linear probes.
Figure 6: Toy setting linking shortcut regions to hallucination detectability. Left: Schematic of the correlation and noisy regions. In the correlation region, shortcut features dominate and induce confident errors that are often missed by detectors. Right: Empirical AUROC of a confidence-based detector versus the proportion of the shortcut region $\varrho$ for a multi-layer perceptron (MLP) and an MLP with only the last layer trained (equivalent to kernel ridgeless regression). Detection performance degrades monotonically as $\varrho$ increases for both models, consistent with the prediction that stronger shortcut reliance yields harder-to-detect hallucinations. Mapping to our toy model: the correlation / shortcut region corresponds to a strongly correlated, low-noise regime; the noisy / idiosyncratic region corresponds to a weakly correlated, high-noise regime.
@@ 5.1 Problem Setup
Data Generation.
Consider a dataset $D_{N}=\left\{(x_{1},y_{1}),\dots,(x_{N},y_{N})\right\}\subset\mathcal{X}\times\mathbb{R}$, where the input space $\mathcal{X}=\mathbb{S}^{d}\subset\mathbb{R}^{d+1}$ is the $d$-dimensional unit sphere. Suppose that $x_{1},\dots,x_{n}$ are drawn i.i.d. from $X\sim\mathrm{Unif}(\mathbb{S}^{d})$, with binary labels $Y\in\left\{+1,-1\right\}$. For a fixed $\varrho\in(0,1)$, the sphere is partitioned into three regions by parallel annuli, resulting in the correlation regions $\mathcal{C}=\mathcal{C}_{+}\cup\mathcal{C}_{-}$ and the noisy region $\mathcal{N}$, such that
$\mathbb{P}\left(X\in\mathcal{C}_{+}\right)=\mathbb{P}\left(X\in\mathcal{C}_{-}\right)=\frac{\varrho}{2}-\frac{\epsilon}{4},\quad\mathbb{P}\left(X\in\mathcal{N}\right)=1-\varrho-\frac{\epsilon}{2},$
where $\epsilon\in\left(0,2\min\left\{\varrho,1-\varrho\right\}\right)$ is introduced to avoid discontinuities in the target function at the boundary separating the correlation and noisy regions, which could otherwise lead to the so-called Gibbs phenomenon (de2020jumping).
@@
Intuitively, the correlation region captures strong but spurious correlations – strong statistical patterns that are not necessarily causal (e.g., persons with surnames ending in kov and Russian birthplaces) – whereas the noisy region exhibits high variance and can only be learned by memorizing individual examples.
Kernel Ridge(less) Regression.
@@
$f_{N,\lambda}(x)=k(x,X_{N})(k(X_{N},X_{N})+N\lambda I_{N})^{-1}Y_{N},$
where $k(x,X_{N})=(k(x,x_{1}),\dots k(x,x_{N}))\in\mathbb{R}^{1\times N}$ is a row vector, $k(X_{N},X_{N})\in\mathbb{R}^{N\times N}$ is the kernel matrix with entries $k(x_{i},x_{j}),1\leq i,j\leq N$, $Y_{N}=(y_{1},\dots,y_{N})^{\mathsf{T}}\in\mathbb{R}^{N}$, and $I_{N}$ denotes the $N\times N$ identity matrix.
When $\lambda=0$, KRR reduces to kernel interpolation, also known as kernel ridgeless regression, which fits all training data points exactly. In this case, the interpolant $f_{N}$ solves the norm-minimizing problem
@@ 5.2 Main Theorem
Hallucination Detection Criterion.
In our setting, prediction values lie in [1,+1][-1,+1], and the model output $f(x)$ can be interpreted as prediction confidence. Hence, a natural approach to detect hallucinations is to impose a fixed threshold $\rho\in(0,1)$ on the output, such that if $|f(x)|<\rho$, the model is deemed to be hallucinating. This criterion relies on the two rules:
@@
    The predictor can reliably distinguish training data points from unseen inputs, i.e., the outputs satisfy $|f(x)|\geq\rho$ for all $x\in X_{N}$.
@@
    The model learns selectively, i.e., it neither ignores all correlations nor indiscriminately learns every correlation.
Our theoretical results show that a broad class of regression models fails to pass confidence-based hallucination detection. Specifically, when the regularization parameter $\lambda>0$, KRR can neither distinguish training data in the noisy region nor detect hallucinations in the correlation region (see Theorem 2 in Appendix B.1), thereby violating rules (i) and (ii). This occurs because the regularization term forces the model to ignore all noisy information while fitting all strong correlations. Conversely, if we reduce the bandwidth to force the model to memorize all data points (see Theorem 7 in Appendix B.2), KRR fails to learn any correlation, thereby violating rule (ii). Therefore, the criterion requires the model to both memorize and generalize.
@@
Under some technical assumptions (see Assumptions 1-4 in Appendix B), let $f_{N}$ be the kernel interpolation solution on the training set $D_{N}$ generated as above. Further, assume either
@@
  • the kernel function is define as $k_{c_{N},\gamma_{N}}(x,x^{\prime})\coloneqq\tilde{k}(x,x^{\prime})+c_{N}\check{k}_{\gamma_{N}}(x,x^{\prime})$, where $\tilde{k}$ is a universal kernel, $\check{k}_{\gamma_{N}}$ is the Laplace kernel with bandwidth $\gamma_{N}>0$, $c_{N}\to 0$, $Nc_{N}^{4}\to\infty$, and $\gamma_{N}\leq N^{-3/d}(7\ln N)^{-1}$.
Then for any $\delta\in(0,1)$, there exist constants $C_{0},N_{0},\alpha>0$ such that for any $N\geq N_{0}$, if we define the uniform upper confidence bound as $U_{N}^{\delta}\coloneqq C_{0}N^{-\alpha}$, the following holds
@@
Therefore, for any threshold $\rho\in(0,0.98)$,
$\liminf_{N\to\infty}\mathbb{P}\left(\mathbb{E}_{D_{N}}\left[|f_{N}(x)|\geq\rho\right]\right)\geq 1-\delta,\quad\text{ for all }x\in\mathcal{C},$
which indicates that any hallucination detection criterion with a fixed $\rho$ fails in the correlation regions (i.e., $f_{N}$ makes confident prediction even for unseen data in $\mathcal{C}$).
For over-parametrized neural networks in the kernel regime, the predictor can be approximated by KRR with neural kernels. When training all layers, the relevant kernel is the neural tangent kernel (NTK) (jacot2018neural); when training only the last layer, it is the neural network Gaussian process (NNGP) kernel (neal1996priors; lee2018deep; matthews2018gaussian). Consequently, Theorem 1 applies to over-parameterized neural networks; see Appendix C for further details on the kernel behavior of neural networks.

### v2

@@ 1 Introduction
Figure 1: Spurious correlations induce high-confidence hallucinations that evade detection and mitigation. Statistical biases in training data (e.g., name-nationality) lead to consistent errors resistant to uncertainty metrics and refusal fine-tuning.
@@ 2.2 Spurious Correlation
Spurious correlations, also called “shortcuts”, are non-causal statistical dependencies in training data and significantly contribute to hallucinations in language models. Recent studies demonstrate that such correlations amplify erroneous outputs, often with high confidence. Multimodal research, for instance, shows object hallucinations being exacerbated by misleading co-occurrences in datasets (hosseini2025seeing; hu2025causal). Similarly, purely textual models suffer from biases like attestation and frequency biases, resulting in incorrect entailments or factual assertions derived from superficial patterns (mckenna2023sources). Conceptual-level spurious correlations are widespread in both fine-tuning and in-context learning settings, posing substantial mitigation challenges across modeling paradigms (zhou2023explore; yuan2024llms). Methods like high-similarity pruning and causal interventions have been proposed to address knowledge-shortcut hallucinations (wang2025kshseek; licausally), yet their efficacy is limited to particular contexts, and their performance under strong correlations remains unclear.
@@
3 Empirical Evaluation of Hallucination Detection and Mitigation under Spurious Correlation
@@ 3.1 Experimental Setting
Figure 2: AUROC of different hallucination detection methods versus $\rho$. Left: Experimental results of pretrained models. Right: Experimental results of models that continue pretrained from SmolLM2-1.7B. The classification performance of different detection methods drops as $\rho$ increases, indicating that spurious correlation hinders hallucination detection.
@@
Following (AllenZhu-icml2024-tutorial; allen2024physics33), we generate profiles for 20,000 individuals, each containing six attributes: date of birth, birth city, university, major, employer, and employer city. To construct both pretraining and supervised fine-tuning datasets, we first design a diverse set of text templates for describing profiles and then embed each individual’s information into natural texts based on these templates (see Table 3 for examples). We uniformly divide the profiles into three subsets—pretraining, instruction fine-tuning, and testing—to ensure balanced representation and prevent overlap. The pretraining set includes the first 10,000 individuals, each represented by 50 diverse text templates; the fine-tuning set uses the first 5,000 of these individuals, generating 30 question–answer pairs per individual. The remaining individuals are reserved exclusively for testing and hallucination detection evaluation. We conduct experiments using GPT2-like models (modded_nanogpt_2024) of various sizes, detailed in Table 4. The training procedures and detailed description of dataset construction are described in Appendix E.
@@
To systematically investigate spurious correlations, we adopt a controlled methodology: each individual’s full name is composed of a first name, middle name, and surname, each randomly selected from distinct sets without repetition. We then associate surnames with specific attributes using a probabilistic mapping that simulates realistic patterns (e.g., surnames ending in kov are often linked to Russian birthplaces). To control correlation strength, we introduce a coefficient $\rho\in[0,1]$, representing the probability that a surname directly determines its associated attribute. With probability $\rho$, the attribute matches the surname-based mapping; otherwise, it is uniformly sampled from all possible values. This approach enables precise manipulation of correlation strength to evaluate existing hallucination detection methods rigorously.
@@ 3.2 Results
We benchmark hallucination detection methods in Table LABEL:tab:hallucination_methods, including perplexity, logit entropy, window entropy, attention score, and linear probing. We selected the linear probing layer that performed best on the training set. As shown in Figure 2, although some methods perform well when $\rho=0$, their performance degrades sharply as $\rho$ increases (e.g., $\rho=0.9$), with most methods failing to maintain reasonable precision.
@@
To verify whether the previously identified spurious-correlation-driven failure persists under a knowledge injection setting, we extend our investigation to real LLMs fine-tuned on synthetic datasets. Using SmolLM2-1.7B (allal2025smollm2) as the base model, we conduct continual pre-training and instruction fine-tuning. As shown in Figure 2, when $\rho$ is high, all evaluated methods still exhibit low precision, confirming that the same failure mode remains even at the 1.7B scale and in the knowledge injection setting.
@@
Figure 3: Performance of fine-tuned models of various sizes under varying correlation coefficients. Left: The test accuracy for factual recall questions regarding known individuals. Right: The refusal rate when queried about unknown individuals.
@@
We investigate how spurious correlations affect refusal fine-tuning, which trains models to reject uncertain or out-of-distribution inputs. Following  zhang2024r; cheng2024can, we add refusal examples during instruction fine-tuning. We keep the fine-tuning format but substitute the entities with unseen individuals, setting the ground truth to I don’t know (see Table 3).
@@
4 Validation on Real World LLM
@@ 4.1 Experimental Setups
We validate our findings on spurious correlations across a diverse set of open-source and commercial models: GPT-5 (openai2025gpt5), DeepSeek V3 (liu2024deepseek), GPT-OSS-20B (agarwal2025gpt), and Qwen3-30B-A3B-Instruct (qwen2025qwen25technicalreport).
@@
For each question–answer pair $(x,y)$, we obtain a consensus model answer $f^{*}(x)$ by running the model $f$ ten times and taking a majority vote over the outputs $f_{1}(x),f_{2}(x),\dots,f_{10}(x)$. We then extract entities from both the question and the consensus answer using an entity extractor $e(\cdot)$ (by prompting LLM) and compute their co-occurrence using the Jaccard similarity (jaccard1908nouvelles):
@@
Intuitively, a higher Jaccard similarity indicates stronger associative priors between the entities in questions and model-generated answers, effectively serving as a proxy for larger $\rho$. We compute these Jaccard similarity scores for all samples and group them into five buckets based on their values, from $T_{1}$ (highest similarity) to $T_{5}$ (lowest). This bucketing allows us to analyze model behavior under different levels of spurious correlation. See case study in Appendix D.1.
@@ 4.2 Results
Figure 5: Hallucination detection performance versus entity co-occurrence. Left: GPT-OSS-20B. Right: Qwen-30B-A3B-Instruct. Classification performance decreases consistently as Jaccard overlap increases, across all evaluated detection methods, including perplexity, window entropy, logit entropy, attention-score heuristics, and linear probes.
@@ 5.1 Problem Setup
Data Generation
Consider a dataset $D_{N}=\left\{(x_{1},y_{1}),\dots,(x_{N},y_{N})\right\}\subset\mathcal{X}\times\mathbb{R}$, where the input space $\mathcal{X}=\mathbb{S}^{d}\subset\mathbb{R}^{d+1}$ is the $d$-dimensional unit sphere. Suppose that $x_{1},\dots,x_{n}$ are drawn i.i.d. from $X\sim\mathrm{Unif}(\mathbb{S}^{d})$, with binary labels $Y\in\left\{+1,-1\right\}$. For a fixed $\rho\in(0,1)$, the sphere is partitioned into three regions (as shown in Figure 6): the correlation regions $\mathcal{C}=\mathcal{C}_{+}\cup\mathcal{C}_{-}$ and the noisy region $\mathcal{N}$, such that
$\mathbb{P}\left(X\in\mathcal{C}_{+}\right)=\mathbb{P}\left(X\in\mathcal{C}_{-}\right)=\frac{\rho}{2}-\frac{\epsilon}{4},\quad\mathbb{P}\left(X\in\mathcal{N}\right)=1-\rho-\frac{\epsilon}{2},$
where $\epsilon\in(0,2\min\{\rho,1-\rho\})$ can be made arbitrarily small. This parameter is introduced to ensure continuity of the target function at region boundaries, thereby mitigating the Gibbs phenomenon (de2020jumping) (further details are provided in Appendix B).
@@
The correlation region captures strong but spurious correlations—statistical patterns that are not necessarily causal (e.g., surnames ending in “kov” and Russian birthplaces)—whereas the noisy region exhibits high variance and can only be learned by memorizing individual examples.
Figure 6: Toy setting linking shortcut regions to hallucination detectability. Left: Schematic of the correlation and noisy regions. In the correlation region, shortcut features dominate and induce confident errors that are often missed by detectors. Right: Empirical AUROC of a confidence-based detector versus the proportion of the shortcut region $\rho$ for a multi-layer perceptron (MLP) and an MLP with only the last layer trained (equivalent to kernel ridgeless regression). Detection performance degrades monotonically as $\rho$ increases for both models, consistent with the prediction that stronger shortcut reliance yields harder-to-detect hallucinations.
Kernel Ridge(less) Regression
@@
$f_{N,\lambda}(x)=k(x,X_{N})(k(X_{N},X_{N})+\lambda NI_{N})^{-1}Y_{N},$
where $k(x,X_{N})=(k(x,x_{1}),\dots k(x,x_{N}))\in\mathbb{R}^{1\times N}$, $k(X_{N},X_{N})=\left(k(x_{i},x_{j})\right)_{1\leq i,j\leq N}\in\mathbb{R}^{N\times N}$ is the kernel matrix, $Y_{N}=(y_{1},\dots,y_{N})^{\mathsf{T}}\in\mathbb{R}^{N}$, and $I_{N}$ denotes the $N\times N$ identity matrix.
For $\lambda=0$, KRR reduces to kernel interpolation, also known as kernel “ridgeless” regression, which interpolates all training data. The resulting interpolant $f_{N}$ solves a norm-minimizing problem:
@@
A key feature of kernel ridgeless regression is its capacity to interpolate training data, closely resembling the behavior exhibited by modern LLMs, where well-trained models must internalize diverse common-sense knowledge. This similarity motivates our focus on kernel regression, offering meaningful insights into how spurious correlations influence LLMs.
@@ 5.2 Main Theorem
We model confidence-based hallucination detection as a binary classification task, and focus on the model’s ability to identify training data and its judgment of spurious correlations.
Hallucination Detection Criterion
Note that the model output $f(x)$ also indicates prediction confidence. An output is classified as a hallucination if its absolute confidence $|f(x)|$ falls below a threshold $\tau\in(0,1)$. This criterion is based on the following two rules:
@@
    The model can reliably distinguish training data from unseen inputs, such that $|f(x)|\geq\tau$ for all $x\in X_{N}$.
@@
    The model exhibits selective learning, avoiding the extremes of either disregarding all spurious correlations or learning them indiscriminately.
Our theoretical results show that a broad class of regression models fails to pass confidence-based hallucination detection. Specifically, when the regularization parameter $\lambda>0$, KRR can neither distinguish training data in the noisy region nor detect hallucinations in the correlation region (see Theorem 2 in Appendix B.1), thereby violating rules (i) and (ii). This happens because the regularization term causes the model to disregard all noisy information while learning all strong correlations. Conversely, if we reduce the bandwidth to make the model memorize all data points (see Theorem 7 in Appendix B.2), KRR fails to learn any correlation, thus violating rule (ii). Therefore, the criterion requires the model to both memorize and generalize.
@@
Under some technical assumptions (see Assumptions 1-4 in Appendix B), let $f_{N}$ be the kernel interpolation solution on the training set $D_{N}$ generated as above. Further, suppose either
@@
  • $k_{c_{N},\gamma_{N}}(x,x^{\prime})\coloneqq\tilde{k}(x,x^{\prime})+c_{N}\check{k}_{\gamma_{N}}(x,x^{\prime})$, where $\tilde{k}$ is a universal kernel, $\check{k}_{\gamma_{N}}$ is the Laplace kernel with bandwidth $\gamma_{N}>0$, $c_{N}\to 0$, $Nc_{N}^{4}\to\infty$, and $\gamma_{N}\leq N^{-3/d}(7\ln N)^{-1}$.
Then for any $\delta\in(0,1)$, there exist constants $C_{0},N_{0},\alpha>0$, for any $N\geq N_{0}$, define the uniform upper confidence bound as $U_{N}^{\delta}\coloneqq C_{0}\delta^{-1}N^{-\alpha}$, the following holds
@@
Therefore, for any threshold $\tau\in(0,0.98)$,
$\liminf_{N\to\infty}\mathbb{P}\left(\mathbb{E}_{D_{N}}\left[|f_{N}(x)|\geq\tau\right]\right)\geq 1-\delta,\quad\text{ for all }x\in\mathcal{C},$
which indicates that any hallucination detection criterion with a fixed $\tau$ fails in the correlation regions (i.e., $f_{N}$ makes confident prediction even for unseen data in $\mathcal{C}$).
In the kernel regime, over-parametrized neural networks can be approximated by kernel ridge regression with neural kernels (see Appendix C for a review). When training all layers, the relevant kernel is the neural tangent kernel (NTK) (jacot2018neural), whereas training only the last layer corresponds to the neural network Gaussian process (NNGP) kernel (neal1996priors; lee2018deep; matthews2018gaussian). Thus, Theorem 1 applies to over-parameterized neural networks, including Transformer architectures in modern LLMs (yang2020neural; yang2021architectural; hron2020infinite). As shown in Figure 6, the increasing spurious correlations consistently impede hallucination detection in fully-connected neural networks, aligning with the findings presented in previous sections.